The combustion process inside a jet engine creates 300 J of heat and does 120 J...

  1. The combustion process inside a jet engine creates 300 J of heat and does 120 J of work spinning a turbine shaft. What is the change in specific enthalpy of the gas?

  1. Through an irreversible process, object A transfers 560 J of heat to object B. The resulting temperature of the system is 241 K. What is the entropy of the system?
  1. For an ideal gas, if the pressure remains the same but the volume is doubled, what would you expect to happen to the temperature? Why?   **Hint: use the equation of state to help justify your answer.
Answer #1

1.

Heat,Q=300 J

Work,W=120 J

Change in specific enthalpy of gas=300-120=180 J

2.

Entropy of system=dQ/T=560/241=2.32 J/K

3.

According to ideal gas equation

PiVi=nRTi ............................1

Pf=Pi

Vf=2Vi

So

Tf=PfVf/nR=Pi2Vi/nR

From 1

Tf=2Ti

So Temperature became doubles.